OpenStudy (anonymous):

I WILL MEDAL + FAN! I WILL MEDAL + FAN! What is the slope of the line? y = -5 - 9 @TheSmartOne @iGreen

OpenStudy (anonymous):

14 cause you subtract 9- (-5) = 14

TheSmartOne (thesmartone):

I think you made a typo...

OpenStudy (anonymous):

@-EtherealEbullience

OpenStudy (anonymous):

wherer is it

hartnn (hartnn):

did u mean y = -5x - 9 ?

TheSmartOne (thesmartone):

Your equation should be in the format of \(y=mx+b\) and m would be your slope...

OpenStudy (gorv):

compare with standard equation of line y=mx+c here m = slope

OpenStudy (anonymous):

Hello, Generally slope-intercept form: \[y = mx + c\] Where \(m\) is the slope and \(c\) is the y-intercept. So I do believe with @TheSmartOne you made a typo... Peace.

OpenStudy (igreen):

Yeah, I think you made a typo. Are you sure its y = -5 - 9? Or is it y = -5x - 9?

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

Yeah, Sorry I made a typo.

OpenStudy (igreen):

@lloio Your answer is correct if she didn't make a typo, you didn't do anything wrong.

OpenStudy (zpupster):

y = -14 x will = 0 so your slop will be 0 the graph will be a straight line at y = -14

OpenStudy (igreen):

Okay, well your equation is in the form of: y = mx + b Where m = slope. So what is the slope? @im.celibate

OpenStudy (anonymous):

yes im been answering

OpenStudy (anonymous):

-4? @iGreen

hartnn (hartnn):

 thats how u compare!